# thanks: Corresponding author

### v1

@@ I Introduction
where, $H_{i}$ is generalized parton distribution (GPD), $A_{i}$ and $D_{i}$ are the gravitational form factors (GFFs), and $i=u,d,g$. The latter point is particularly interesting because it provides access to the individual quark and gluon distributions. Formally, the individual GFFs can be obtained from the Lorentz decomposition of the hadronic matrix element (HME). For a scalar, such as the pion, the decomposition reads Polyakov:2018zvc ,
@@
where, $P=\frac{1}{2}(p^{\prime}+p)$, and $q=p^{\prime}-p$, $Q^{2}=-q^{2}$. $A_{i}(-q^{2},\mu)$, $D_{i}(-q^{2},\mu)$ and $\bar{c}_{i}(-q^{2},\mu)$ are the GFFs as mentioned. $\mu$ represents the renormalization scale, i.e. the hadronic EMT are renormalization scheme and scale dependent. Unless elsewhere stated, we will suppress the dependence on $\mu$ hereafter for simplicity.
@@
where, $e_{i}$, $p_{i}$ and $\pi^{\alpha\beta}_{i}$ are the physical densities associated with an individual constituent. A new density $\Lambda_{i}$, in analogy to cosmological constant Teryaev:2013qba ; Teryaev:2016edw ; Liu:2023cse , emerges due to the non-conserving nature of the individual EMT $\mathcal{T}^{\alpha\beta}_{i}$. It represents the external pressure acting on the $i$-th constituent, and is related to $\bar{c}_{i}$ by Fourier transform. The physical densities defined in (3) are the sum of the individual densities, e.g., the energy density, $e=\sum_{i}e_{i}$. In hydrodynamics, such a decomposition is known as the coupled multi-fluid picture Rezzolla:2013dea . In this picture, all fluid species have the same four-velocity, which is in contrast to the interacting multi-fluid picture. The latter is more suitable for describing loosely bound hadrons, e.g. hadron molecules and nuclei He:2024jgc .
@@
In this work, we investigate the decomposition of the hadronic EMT within a (3+1)D strongly coupled scalar theory. This is a continuation of our previous work on the evaluation of the GFFs from the same theory Cao:2023ohj . The simplicity of the scalar theory allows us to apply the light-front Hamiltonian approach with a systematic Fock space expansion even in the strong-coupling regime Li:2014kfa ; Li:2015iaw ; Karmanov:2016yzu . The method provides access to the microscopic description of the GFFs as well as the related physical densities with a wave function representation, in analogy to the famous Drell-Yan-West formula for the electromagnetic form factors Drell:1969km ; West:1970av ; Brodsky:1980zm ; Brodsky:1998hn ; Brodsky:2000ii . One of the main challenges to compute the GFFs within the light-front Hamiltonian formalism is the non-perturbative renormalization of the energy-momentum tensor operator. Since the light-front formulation is not manifestly covariant, the divergences of different components of the operator could be dramatically different. To extract GFFs $A(Q^{2})$ and $D(Q^{2})$, we employed operators $T^{++}$ and $T^{+-}$ Cao:2023ohj . Here, we adopt light-front coordinates, $x^{\pm}=x^{0}+x^{3}$ and $\vec{x}_{\perp}=(x^{1},x^{2})$. In particular, we showed that $T^{+-}$ is properly renormalized using the counterterms for the light-front Hamiltonian $P^{-}$. Indeed, $T^{+-}$ is the density of the light-front Hamiltonian Brodsky:1997de . To access the individual GFFs $A_{i}(Q^{2})$, $D_{i}(Q^{2})$ and $\bar{c}_{i}(Q^{2})$, we need one more current. One of the natural choices is $T^{11}_{i}$ (or $T^{22}_{i}$ or their sum); however one can show that the HME is divergent, even for the total current $T^{11}$. Note that using the divergence of the $T^{\mu 1}_{i}$ is equivalent to adopting $T^{11}_{i}$. In Ref. Cao:2024rul , we performed a systematic covariant analysis of the HME of the light-front EMT based on a comparison of the covariant and light-front perturbation theory. We found that $T^{11}_{i}$ and $T^{22}_{i}$ are contaminated by the so-called spurious form factors, which contain uncanceled UV divergences. We also concluded that $T_{i}^{++}$, $T^{+-}_{i}$ and $T_{i}^{12}$ are “good currents” that are free of the spurious form factors and can be used to extract the full set of GFFs. In this work, we apply this formulation to the scalar Yukawa theory in the non-perturbative regime.
@@ IV Numerical results
The non-vanishing $\bar{c}(Q^{2})$ represents an isotropic spurious external force distribution on the transverse plane. For the scalar Yukawa model, the force is attractive at the center $r_{\perp}=0$ and repulsive elsewhere – the pion cloud is slightly stretched outward. Since this spurious force originates from the Fock space truncation, which implicitly breaks the Poincaré symmetry, we expect that incorporating higher Fock sector contributions would reduce its magnitude Zhang:2024 , which is already very small in the present work.
@@ Acknowledgements
The authors acknowledge fruitful discussions with S. Xu, C. Mondal, S. Nair, X. Zhao and V.A. Karmanov. Y.L. is supported by the National Natural Science Foundation of China (NSFC) under Grant No. 12375081, by the Chinese Academy of Sciences under Grant No. YSBR-101. J.P.V. acknowledges support from the US Department of Energy (DOE) under Grant No. DE-SC0023692.

### v2

@@ I Introduction
where, $H_{i}$ is generalized parton distribution (GPD), $A_{i}$ and $D_{i}$ are the gravitational form factors (GFFs), and $i=u,d,g$ refers to each particle species. The latter point is particularly interesting because it provides access to the individual quark and gluon distributions. Formally, the individual GFFs can be obtained from the Lorentz decomposition of the hadronic matrix element (HME). For a scalar, such as the pion, the decomposition reads Polyakov:2018zvc ,
@@
where, $P=\frac{1}{2}(p^{\prime}+p)$, and $q=p^{\prime}-p$, $Q^{2}=-q^{2}$. $A_{i}(-q^{2},\mu)$, $D_{i}(-q^{2},\mu)$ and $\bar{c}_{i}(-q^{2},\mu)$ are the GFFs as mentioned. $i$ refers to specific particle species, e.g. quark. $\mu$ represents the renormalization scale, i.e. the hadronic EMT are renormalization scheme and scale dependent. Unless elsewhere stated, we will suppress the dependence on $\mu$ hereafter for simplicity.
@@
where, $e_{i}$, $p_{i}$ and $\pi^{\alpha\beta}_{i}$ are the physical densities associated with an individual constituent. A new density $\Lambda_{i}$, in analogy to cosmological constant Teryaev:2013qba ; Teryaev:2016edw ; Liu:2023cse , emerges due to the non-conserving nature of the individual EMT $\mathcal{T}^{\alpha\beta}_{i}$, i.e. EMT of each particle species. It represents the external pressure acting on the $i$-th constituent, and is related to $\bar{c}_{i}$ by Fourier transform. The physical densities defined in (3) are the sum of the individual densities, e.g., the energy density, $e=\sum_{i}e_{i}$. In hydrodynamics, such a decomposition is known as the coupled multi-fluid picture Rezzolla:2013dea . In this picture, all fluid species have the same four-velocity, which is in contrast to the interacting multi-fluid picture. The latter is more suitable for describing loosely bound hadrons, e.g. hadron molecules and nuclei He:2024jgc .
@@
In this work, we investigate the decomposition of the hadronic EMT within a (3+1)D strongly coupled scalar theory. This is a continuation of our previous work on the evaluation of the GFFs from the same theory Cao:2023ohj . The simplicity of the scalar theory allows us to apply the light-front Hamiltonian approach with a systematic Fock space expansion even in the strong-coupling regime Li:2014kfa ; Li:2015iaw ; Karmanov:2016yzu . The method provides access to the microscopic description of the GFFs as well as the related physical densities with a wave function representation, in analogy to the famous Drell-Yan-West formula for the electromagnetic form factors Drell:1969km ; West:1970av ; Brodsky:1980zm ; Brodsky:1998hn ; Brodsky:2000ii . One of the main challenges to compute the GFFs within the light-front Hamiltonian formalism is the non-perturbative renormalization of the energy-momentum tensor operator. Since the light-front formulation is not manifestly covariant, the divergences of different components of the operator could be dramatically different. To extract GFFs $A(Q^{2})$ and $D(Q^{2})$, we employed operators $T^{++}$ and $T^{+-}$ Cao:2023ohj . Here, we adopt light-front coordinates, $x^{\pm}=x^{0}\pm x^{3}$ and $\vec{x}_{\perp}=(x^{1},x^{2})$. In particular, we showed that $T^{+-}$ is properly renormalized using the counterterms for the light-front Hamiltonian $P^{-}$. Indeed, $T^{+-}$ is the density of the light-front Hamiltonian Brodsky:1997de . To access the individual GFFs $A_{i}(Q^{2})$, $D_{i}(Q^{2})$ and $\bar{c}_{i}(Q^{2})$, we need one more current. One of the natural choices is $T^{11}_{i}$ (or $T^{22}_{i}$ or their sum); however one can show that the HME is divergent, even for the total current $T^{11}$. Note that using the divergence of the $T^{\mu 1}_{i}$ is equivalent to adopting $T^{11}_{i}$. In Ref. Cao:2024rul , we performed a systematic covariant analysis of the HME of the light-front EMT based on a comparison of the covariant and light-front perturbation theory. We found that $T^{11}_{i}$ and $T^{22}_{i}$ are contaminated by the so-called spurious form factors, which contain uncanceled UV divergences. We also concluded that $T_{i}^{++}$, $T^{+-}_{i}$ and $T_{i}^{12}$ are “good currents” that are free of the spurious form factors and can be used to extract the full set of GFFs. In this work, we apply this formulation to the scalar Yukawa theory in the non-perturbative regime.
@@ IV Numerical results
The non-vanishing $\bar{c}(Q^{2})$ represents an isotropic spurious external force distribution on the transverse plane. Note that because $\bar{c}(0)=0$, the light-front energy is still conserved after integrating the light front energy density $T^{+-}$ over the entire space. For the scalar Yukawa model, the force is attractive at the center $r_{\perp}=0$ and repulsive elsewhere – the pion cloud is slightly stretched outward. Since this spurious force originates from the Fock space truncation, which implicitly breaks the Poincaré symmetry, we expect that incorporating higher Fock sector contributions would reduce its magnitude Zhang:2025wli , which is already very small in the present work.
@@ Acknowledgements
The authors acknowledge fruitful discussions with S. Xu, C. Mondal, S. Nair, X. Zhao and V.A. Karmanov. Y.L. is supported by the National Natural Science Foundation of China (NSFC) under Grant No. 12375081, by the Chinese Academy of Sciences under Grant No. YSBR-101.